# Vector meson’s spin alignments in high energy reactions

### v1

@@ Abstract
The global spin alignment of vector mesons has been observed by the STAR collaboration at the Relativistic Heavy Ion Collider (RHIC) at Brookhaven National Laboratory (BNL). It provides a unique opportunity to probe the correlation between the polarized quark and antiquark in the strongly coupled quark-gluon plasma (sQGP) produced in relativistic heavy ion collisions, opening a new window to explore the properties of sQGP. In addition, spin alignments of vector mesons have also been observed in other high-energy particle collisions. The results seem to be strongly dependent on the hadronization mechanism, so comprehensive studies are needed. In this article, we present a brief review of theoretical and experimental advances in the study of vector meson’s spin alignments in a variety of high-energy particle collisions, with emphasis on hadronization mechanisms.
@@ I Introduction
Quantum Chromodynamics (QCD) is the fundamental theory for strong interaction that is responsible for binding quarks and gluons together to form protons, neutrons, and other hadrons. Relativistic Heavy Ion Collisions (RHIC) are a powerful tool for studying Quantum Chromodynamics (QCD) and properties of strongly interacting matter under extreme conditions Braun-Munzinger:2015hba ; Florkowski:2018fap ; Ko ; Rapp ; Zhang1 ; ZhaoJ . In these collisions, heavy nuclei are accelerated to nearly light speed and then collide to achieve immense energy density that recreates conditions similar to those just microseconds after the Big Bang, allowing scientists to study quark-gluon plasma (QGP) - a state of matter in which quarks and gluons are decoupled from hadrons Pandav:2022xxx ; HeWB ; Chen:2023mel ; MaYG ; PangLG ; Sun1 ; Chen1 .
Spin is an intrinsic form of angular momentum carried by elementary particles, and it is a (pseudo-)vector quantity that can point to different directions. As elementary particles of strong interaction, quarks and gluons carry spins. As composite particles of strong interaction, many hadrons also carry spins. The spin polarization refers to the alignment of the spin along a specific direction (called spin quantization direction). Spin degrees of freedom have been playing an important role in the development of modern physics since its discovery in 1925. The global polarization effect (GPE) of the quark gluon plasma (QGP) produced in high energy heavy-ion collisions is a new spin effect in particle and nuclear physics. The theoretical prediction Liang:2004ph ; Liang:2004xn was made almost two decades ago and attracted immediate attention Voloshin:2004ha ; Betz:2007kg ; Ipp:2007ng ; Becattini:2007sr ; Liang:2007ma ; Chen:2007zzq ; Gao:2007bc ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu . However the enthusiasm was soon dampened by STAR’s earlier attempts Chen:2007zzq ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu that gave null results for global $\Lambda$-hyperon polarization as well as global spin alignments of vector mesons due to limited statistics in data at that time. The earlier results show that the GPE, even if it exists, could be very small, so it is unclear if it is within the scope and resolution of current experiments. The enthusiasm was aroused by non-vanishing results for the GPE for $\Lambda$ hyperons in STAR’s beam-energy-scan experiments STAR:2017ckg , which showed that the GPE decreases monotonically with collision energies. Now the GPE has grown to be a sub-field in heavy-ion collisions, see Refs. Wang:2017jpl ; Liang:2019clf ; Gao:2020lxh ; Huang:2020dtn ; Gao:2020vbh ; Becattini:2020ngo ; Becattini:2022zvf ; Becattini:2024uha for recent reviews.
@@
In addition to relativistic heavy-ion collisions, the vector meson’s spin alignment has also been studied in other high energy processes such as $e^{+}e^{-}$, $e^{-}p$ and $pp$ collisions DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs ; NOMAD:2006kuc . The results show quite different features and theoretical efforts have also been made to describe them Anselmino:1997ui ; Anselmino:1998jv ; Anselmino:1999cg ; Xu:2001hz ; Xu:2003fq ; Chen:2016moq ; Chen:2016iey ; Chen:2020pty . It is therefore desirable to summarize these experimental and theoretical results and make comparison among different high energy reactions as guidance for future studies.
@@ II.1 Spin density matrices
where the spin quantization direction is assumed to be $\hat{z}$ direction.
@@ II.2 Measurement of spin density matrix through angular distribution of decay product
where $\rho^{A}_{m_{A}m^{\prime}_{A}}=\langle m_{A}|\hat{\rho}^{A}|m^{\prime}_{A}\rangle$ denotes elements of $\hat{\rho}^{A}$.
@@
(1) $j_{A}=1/2$, $\lambda_{1}=\pm 1/2$, $\lambda_{2}=0$, as in a spin-1/2 baryon’s decay into a spin-1/2 baryon and a pion, $H\to B\pi$. In this case, we have two independent helicity amplitudes $H_{A}(\pm 1/2,0)$ and $W(\theta,\varphi)$ is given by
@@
Measurements on hyperon polarization and vector meson’s spin alignment have been carried out in different high energy reactions using Eqs. (30-34) (see e.g. Refs. Lesnik:1975my ; Bunce:1976yb ; Bensinger:1983vc ; Gourlay:1986mf ; TASSO:1984nda ; ALEPH:1996oew ; OPAL:1997oem , DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs ; NOMAD:2006kuc and Chen:2007zzq ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu ). In this brief review, we concentrate on spin alignment of vector mesons with light flavors in comparison with hyperon polarization, thus only Eqs. (30) and (32) are involved.
@@ III.1 Vector meson’s spin alignments in $e^{+}e^{-}$-annihilations
The earliest measurements of vector meson’s spin alignments in high energy reactions might be made at the Large Electron-Position collider (LEP) at European Organization for Nuclear Research (CERN) in the 1990s DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs , where one of the popular polarization axis was the helicity axis defined by the individual particle’s momentum direction. Measurements have been carried out by DELPHI and OPAL Collaborations for $K^{*0}$ and $\phi$ mesons and even for heavy flavor meson such $B$ and $D^{*}$ DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs . They also measured the off-diagonal element such as $\rho_{1,-1}$. As an example, we show the results from Ref. OPAL:1997vmw in Fig. 1.
@@ III.2 Global spin alignments of vector mesons in heavy-ion collisions
Measurements of the vector meson’s spin alignment in heavy-ion collisions are different from conventional studies. A new polarization axis along the nucleus-nucleus system’s orbital angular momentum can be defined Liang:2004ph ; Liang:2004xn .
@@
According to various studies, there are many sources that contribute to the global spin alignment of $\phi$ mesons including vortical flows Yang:2017sdk , electromagnetic fields Yang:2017sdk generated by the electric currents carried by the colliding nuclei, local spin alignment Xia:2020tyd , quark polarization along the direction of its momentum (helicity polarization) Gao:2021rom , the spin alignment by fragmentation Liang:2004xn of polarized quarks, and the shear stress tensor Li:2022vmb ; Wagner:2022gza ; Dong:2023cng . However, these conventional mechanisms are not sufficient to account for the observed $\rho_{00}$ for $\phi$ mesons. It was also proposed that local correlations or fluctuations in turbulent color fields Muller:2021hpe and glasma fields Kumar:2023ghs can also generate a significant contribution to $\rho_{00}$. A recent theoretical development based on local correlations or fluctuations of $\phi$ vector fields can describe the experimental data. The $\phi$ vector field is the ’33’ component of the SU(3) vector multiplet induced by currents of pseudo-Goldstone bosons Manohar:1983md , and it can polarize $s$ and $\bar{s}$ quarks in the same way as the electromagnetic field does. The solid curve of Fig. 3 is such a fit to the data. We see that the calculation describes the data reasonably well. The correlation of the $\phi$ vector field can be quantified by $\langle{\vec{E}}_{\phi}^{2}\rangle$ and $\langle{\vec{B}}_{\phi}^{2}\rangle$ and can be extracted by fitting data, where ${\vec{E}}_{\phi}$ and ${\vec{B}}_{\phi}$ are electric and magnetic components of the $\phi$ vector field respectively. We will discuss more details about the theoretical model in Sect. IV.2.
@@ IV.1 Vector meson’s spin alignments in non-relativistic quark combination models
From Eqs. (46) and (47), we see that, in contrast to $P_{H}$, $\rho^{V}_{00}$ is quadratic in $P_{q}$ and should be less than $1/3$. We emphasize that the case considered in Ref. Liang:2004xn is over simplified in the fact that only the spin degree of freedom for the quark and antiquark is considered, neglecting other degree(s) of freedom and the correlation among the quark’s and antiquark’s polarization. So it might not be a surprise that the STAR data STAR:2017ckg ; STAR:2022fan for the global spin alignment of $\phi$ mesons show a large deviation from 1/3, far beyond the value estimated from $P_{\Lambda}$’s data by its square. In fact, if we make a step forward by considering the dependence of $P_{q}$ on other degree of freedom Sheng:2019kmk ; Sheng:2020ghv ; Sheng:2022ffb ; Sheng:2022wsy $P_{q_{1}}P_{\bar{q}_{2}}$ in Eq. (46) should be replaced by $\langle P_{q_{1}}P_{\bar{q}_{2}}\rangle$, so we have
@@ IV.2.2 Relativistic model in quantum kinetic approach
where $I_{B,i}({\vec{p}})$ and $I_{E,i}({\vec{p}})$ are momentum functions. One can further take the space and momentum average of $\rho_{00}^{\phi}(x,{\vec{p}})$
@@ VII Summary and outlook
It is impressive and interesting to see the distinction between the global spin alignments in high energy heavy-ion collisions and in $e^{+}e^{-}$ or $pp$ collisions which are dominated by different hadronization mechanisms. It is important to extend the measurements to different high energy reactions at different energies to test the universality of these properties. Obviously such studies provide new insights into properties of sQGP and hadronization mechanisms.
@@ Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under Contract No. 2022YFA1604900, by the National Natural Science Foundation of China (NSFC) under Contract Nos. 12025501, 11890710, 11890713, 11890714, 12147101 and 12135011, and by the Strategic Priority Research Program of the Chinese Academy of Sciences (CAS) under Grant No. XDB34030102.

### v2

@@ Abstract
The global spin alignment of vector mesons has been observed by the STAR collaboration at the Relativistic Heavy Ion Collider (RHIC) at Brookhaven National Laboratory (BNL). It provides a unique opportunity to probe the correlation between the polarized quark and antiquark in the strongly coupled quark-gluon plasma (sQGP) produced in relativistic heavy ion collisions, opening a new window to explore the properties of sQGP. In addition, spin alignments of vector mesons have also been observed in other high-energy particle collisions such as $e^{+}e^{-}$ annihilations at high energies where hadron production is dominated by quark fragmentation mechanism. The results obtained are quite different from those obtained in heavy ion collisions where quark coalescence/combination mechanism dominates suggesting strong dependence on hadronization mechanisms. So comprehensive studies in different hadronization processes are needed. In this article, we present a brief review of theoretical and experimental advances in the study of vector meson’s spin alignments in a variety of high-energy particle collisions, with emphasis on hadronization mechanisms.
Keywords: Vector meson, spin alignment, global polarization, relativistic heavy ion collisions, fragmentation
PACS: 12.38.Mh,12.38.–t, 13.88.+e, 13.66.Bc, 25.75.–q, 25.75.Nq
@@ I Introduction
Quantum Chromodynamics (QCD) is the fundamental theory for strong interaction that is responsible for binding quarks and gluons together to form protons, neutrons, and other hadrons. Relativistic Heavy Ion Collisions (RHIC) are a powerful tool for studying QCD and properties of strongly interacting matter under extreme conditions Braun-Munzinger:2015hba ; Florkowski:2018fap ; Ko ; Rapp ; Zhang1 ; ZhaoJ . In these collisions, heavy nuclei are accelerated to nearly light speed and then collide to achieve immense energy density that recreates conditions similar to those just microseconds after the Big Bang, allowing scientists to study quark-gluon plasma (QGP) - a state of matter in which quarks and gluons are decoupled from hadrons Pandav:2022xxx ; HeWB ; Chen:2023mel ; MaYG ; PangLG ; Sun1 ; Chen1 .
Spin is an intrinsic form of angular momentum carried by elementary particles, and it is a (pseudo-)vector quantity that can point to different directions. As elementary particles of strong interaction, quarks and gluons carry spins. As composite particles of strong interaction, many hadrons also carry spins. The spin polarization refers to the alignment of the spin along a specific direction (called spin quantization direction). Spin degrees of freedom have been playing an important role in the development of modern physics since its discovery in 1925. The global polarization effect (GPE) of the quark gluon plasma (QGP) produced in high energy heavy-ion collisions is a new spin effect in particle and nuclear physics. The theoretical prediction Liang:2004ph ; Liang:2004xn ; Gao:2007bc was made almost two decades ago and attracted immediate attention Voloshin:2004ha ; Betz:2007kg ; Ipp:2007ng ; Becattini:2007sr ; Liang:2007ma ; Chen:2007zzq ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu . However the enthusiasm was soon dampened by STAR’s earlier attempts Chen:2007zzq ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu that gave null results for global $\Lambda$-hyperon polarization as well as global spin alignments of vector mesons due to limited statistics in data at that time. The earlier results show that the GPE, even if it exists, could be very small, so it is unclear if it is within the scope and resolution of current experiments. The enthusiasm was aroused by non-vanishing results for the GPE for $\Lambda$ hyperons in STAR’s beam-energy-scan experiments STAR:2017ckg , which showed that the GPE decreases monotonically with collision energies. Now the GPE has grown to be a sub-field in heavy-ion collisions, see Refs. Wang:2017jpl ; Liang:2019clf ; Gao:2020lxh ; Huang:2020dtn ; Gao:2020vbh ; Becattini:2020ngo ; Becattini:2022zvf ; Becattini:2024uha for recent reviews.
@@
In addition to relativistic heavy-ion collisions, the vector meson’s spin alignment has also been studied in other high energy processes such as $e^{+}e^{-}$, $e^{-}p$ and $pp$ collisions DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs ; ZEUS:1999wqh ; NOMAD:2006kuc ; STAR:2008lcm . The results show quite different features and theoretical efforts have also been made to describe them Anselmino:1997ui ; Anselmino:1998jv ; Anselmino:1999cg ; Xu:2001hz ; Xu:2003fq ; Chen:2016moq ; Chen:2016iey ; Chen:2020pty . It is therefore desirable to summarize these experimental and theoretical results and make comparison among different high energy reactions as guidance for future studies.
@@ II.1 Spin density matrices
where the spin quantization direction is assumed to be $\hat{z}$ direction. For the measurement of global spin alignment, the spin quantization direction is taken as the direction of global orbital angular momentum, i.e., $\vec{\varepsilon}_{m=0}=\hat{y}$, while other polarization vectors are determined by conditions $\vec{\varepsilon}_{m}^{\,*}\cdot\vec{\varepsilon}_{m^{\prime}}=\delta_{mm^{\prime}}$ and $\vec{\varepsilon}_{m=\pm 1}^{\,*}=-\vec{\varepsilon}_{m=\mp 1}$. Another widely used choice is the direction of meson’s three-momentum in the laboratory frame, $\vec{\varepsilon}_{m=0}=\vec{p}/|\vec{p}|$, corresponding to the helicity polarization. The spin density matrix for different choices of spin quantization directions are related by the following transformation in the spin space,
$\hat{\rho}_{V}=R\,\hat{\rho}^{\prime}_{V}\,R^{\dagger},$ (22)
where $\hat{\rho}_{V}$ and $\hat{\rho}^{\prime}_{V}$ are two spin density matrices corresponding to two set of polarization vectors $\vec{\varepsilon}_{m=0,\pm 1}$ and $\vec{\varepsilon}^{\,\prime}_{m=0,\pm 1}$, respectively. The transformation matrix $R$ is a $3\times 3$ matrix in the spin space with elements given by the inner product of polarization vectors, $R_{mm^{\prime}}\equiv\vec{\varepsilon}^{\,*}_{m}\cdot\vec{\varepsilon}^{\,\prime}_{m^{\prime}}$.
@@ II.2 Measurement of spin density matrix through angular distribution of decay product
where $\rho^{A}_{m_{A}m^{\prime}_{A}}=\langle m_{A}|\hat{\rho}^{A}|m^{\prime}_{A}\rangle$ denotes elements of $\hat{\rho}^{A}$ and $N$ is the normalization constant.
@@
(1) $j_{A}=1/2$, $\lambda_{1}=\pm 1/2$, $\lambda_{2}=0$, as in a spin-1/2 hyperon’s decay into a spin-1/2 baryon and a pion, $H\to B\pi$, where $H$ denotes a hyperon. In this case, we have two independent helicity amplitudes $H_{A}(\pm 1/2,0)$ and $W(\theta,\varphi)$ is given by
@@
Measurements on hyperon polarization and vector meson’s spin alignment have been carried out in different high energy reactions using Eqs. (31-35) (see e.g. Refs. Lesnik:1975my ; Bunce:1976yb ; Bensinger:1983vc ; Gourlay:1986mf ; TASSO:1984nda ; ALEPH:1996oew ; OPAL:1997oem , DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs ; NOMAD:2006kuc and Chen:2007zzq ; Selyuzhenkov:2007ab ; STAR:2008lcm ; STAR:2007ccu ). In this brief review, we concentrate on spin alignment of vector mesons with light flavors in comparison with hyperon polarization, thus only Eqs. (31) and (33) are involved.
@@ III.1 Vector meson’s spin alignments in $e^{+}e^{-}$-annihilations
The earliest measurements of vector meson’s spin alignments in high energy reactions might be made at the Large Electron-Position collider (LEP) at European Organization for Nuclear Research (CERN) in the 1990s DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs , where one of the popular polarization axis was the helicity axis defined by the individual particle’s momentum direction. Measurements have been carried out by DELPHI and OPAL Collaborations for $K^{*0}$ and $\phi$ mesons and even for heavy flavor meson such as $B$ and $D^{*}$ DELPHI:1997ruo ; OPAL:1997vmw ; OPAL:1997nwj ; OPAL:1999hxs . They also measured the off-diagonal element such as $\rho_{1,-1}$. As an example, we show the results from Ref. OPAL:1997vmw in Fig. 1.
@@ III.2 Global spin alignments of vector mesons in heavy-ion collisions
New measurements of the vector meson’s spin alignment in heavy-ion collisions are different from conventional studies, where a new polarization axis along the nucleus-nucleus system’s orbital angular momentum is defined Liang:2004ph ; Liang:2004xn , and this is the so-called global spin alignment. In experiment, the quantization axis is determined by the normal of the reaction plane, which can be reconstructed by using the charge particle momentum distribution collected in the detector STAR:2022fan . Particles of interest, for example, the vector-meson $\phi$ and $K^{*0}$ are observed by paring of their decay daughters ($K^{\pm}$, $\pi^{\pm}$) with subtraction of the combinatorial background. Then the polar angle distribution of Eq. (34) is analyzed, and the $\rho_{00}$ is extracted after correction for detection efficiency and acceptance STAR:2022fan .
@@
According to various studies, there are many sources that contribute to the global spin alignment of $\phi$ mesons including vortical flows Yang:2017sdk , electromagnetic fields Yang:2017sdk generated by the electric currents carried by the colliding nuclei, local spin alignment Xia:2020tyd , quark polarization along the direction of its momentum (helicity polarization) Gao:2021rom , the spin alignment by fragmentation Liang:2004xn of polarized quarks, and the shear stress tensor Li:2022vmb ; Wagner:2022gza ; Dong:2023cng . However, these conventional mechanisms are not sufficient to account for the observed $\rho_{00}$ for $\phi$ mesons. It was also proposed that local correlations or fluctuations in turbulent color fields Muller:2021hpe and glasma fields Kumar:2023ghs can also generate a significant contribution to $\rho_{00}$. A recent theoretical development based on local correlations or fluctuations of $\phi$ vector fields can describe the experimental data. The $\phi$ vector field is the ’33’ component of the SU(3) vector multiplet induced by currents of pseudo-Goldstone bosons Manohar:1983md , and it can polarize $s$ and $\bar{s}$ quarks in the same way as the electromagnetic field does. The solid curve of Fig. 3 is such a fit to the data. We see that the calculation describes the data reasonably well. The correlation of the $\phi$ vector field can be quantified by $\langle{\vec{E}}_{\phi}^{2}\rangle$ and $\langle{\vec{B}}_{\phi}^{2}\rangle$ and can be extracted by fitting data, where ${\vec{E}}_{\phi}$ and ${\vec{B}}_{\phi}$ are electric and magnetic components of the $\phi$ vector field respectively. We will discuss more details about the theoretical model in Sec. IV.2.
@@ IV.1 Vector meson’s spin alignments in non-relativistic quark combination models
From Eqs. (47) and (48), we see that, in contrast to $P_{H}$, $\rho^{V}_{00}$ is quadratic in $P_{q}$ and should be less than $1/3$. We emphasize that the case considered in Ref. Liang:2004xn is over simplified in the fact that only the spin degree of freedom for the quark and antiquark is considered, neglecting other degree(s) of freedom and the correlation among the quark’s and antiquark’s polarization. So it might not be a surprise that the STAR data STAR:2017ckg ; STAR:2022fan for the global spin alignment of $\phi$ mesons show a large deviation from 1/3, far beyond the value estimated from $P_{\Lambda}$’s data by its square. In fact, if we make a step forward by considering the dependence of $P_{q}$ on other degree of freedom Sheng:2019kmk ; Sheng:2020ghv ; Sheng:2022ffb ; Sheng:2022wsy , $P_{q_{1}}P_{\bar{q}_{2}}$ in Eq. (47) should be replaced by $\langle P_{q_{1}}P_{\bar{q}_{2}}\rangle$, so we have
@@ IV.2.2 Relativistic model in quantum kinetic approach
where $I_{B,i}({\vec{p}})$ and $I_{E,i}({\vec{p}})$ are momentum functions. These functions are derived by substituting transformations (IV.2.2) into Eq. (63) and then separating different field components. For the global spin alignment, the explicit expression of $I_{B,i}$ reads Sheng:2022ffb
$I_{B,i}(p)$ $=$ $C_{1}\left[(E_{p}^{\phi})^{2}+\left(\frac{6E_{p}^{\phi}(p_{2})^{2}}{m_{\phi}+E_{p}^{\phi}}-3(E_{p}^{\phi})^{2}\right)\delta_{2,i}\right.$ (67)
$\left.-\left(1+\frac{3(p_{2})^{2}}{(m_{\phi}+E_{p}^{\phi})^{2}}\right)(p_{i})^{2}\right]$
$+C_{2}\left[\vec{p}\cdot\vec{p}-(p_{i})^{2}-3\left(\sum_{j=1,2,3}\epsilon_{2ij}p_{j}\right)^{2}\right],$
where $\epsilon$ in the last line is the rank-3 antisymmetric tensor. The explicit expression of $I_{E,i}(p)$ can be obtained from $I_{B,i}$ by exchanging $C_{1}$ with $C_{2}$. One can further take the space and momentum average of $\rho_{00}^{\phi}(x,{\vec{p}})$
@@ VII Summary and outlook
It is impressive and interesting to see the distinction between the global spin alignments in high energy heavy-ion collisions and in $e^{+}e^{-}$ or $pp$ collisions which are dominated by different hadronization mechanisms. It is important to extend the measurements to different high energy reactions at different energies to test the universality of these properties. Further measurements on global spin alignments of different vector mesons in relativistic heavy ion collisions are expected to be carried out in the near future by the STAR Collaboration at RHIC and ALICE Collaboration at LHC. More studies are also planned at GSI, HIAF and NICA in the lower energy regions. Also studies in $pp$ collisions are expected by STAR at RHIC, and in $e^{+}e^{-}$ by BESIII at BEPC and Belle II at KEK. They can also be measured in lepton-nucleon scatterings at future EIC. Obviously such studies provide new insights into properties of sQGP and hadronization mechanisms.
@@ Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under Contract No. 2022YFA1604900, by the National Natural Science Foundation of China (NSFC) under Contract Nos. 12025501, 11890710, 11890713, 11890714, 12147101 and 12135011, by the Strategic Priority Research Program of the Chinese Academy of Sciences (CAS) under Grant No. XDB34030102 and by the Natural Science Foundation of Shandong Province.